Problem 3 ABC Bank offers an interest rate of 6.5% on both savings and loans. On the other hand, XYZ Bank offers an interestplease answer with explanation
Answer #1

a. ABC Bank has the lower interest rate of 6.5% on both savings and loans. XYZ Bank offers a higher rate of 7% on both savings and loans. So, the opportunity exists.

Borrow from ABC Bank at a lower interest rate of 6.5% and deposit the same amount in XYZ Bank and earn 7% on the deposit amount.

Example: Let's borrow $100,000 from ABC Bank and deposit the amount in XYZ bank for 1 year.

After one year we owe 100,000 * (1 + 0.065) = $106,500 to ABC Bank.

Our deposit amount in ZYZ Bank would have grown to 100,000 * (1 + 0.07) = $107,000

We withdraw the deposit amount of 107,000 and payback our loan of 106,500. We are left with 107,000 - 106,500 = $500 profit for one year. An arbitrage profit of 500/100,000 = 0.5%.

b. ABC Bank will experience a surge in demand for loans because it offers a lower interest rate of 6.5% as compared to the loan from XYZ Bank at 7%.

XYZ Bank will experience a surge in demand for deposits because it offers a higher interest rate of 7% on the deposits as compared to the ABC Bank at 6.5%.

Can you please upvote? Thank You :-)